OpenStudy (anonymous):

Compute the indefinite integral. Calc 2 help!

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ x+1 }{ \sqrt{x^2+2x+7} } dx\]

OpenStudy (anonymous):

Okay, so I am a bit lost? lol

OpenStudy (isaiah.feynman):

Don't worry Albert Einstein was lost too.

OpenStudy (anonymous):

haha true :P

OpenStudy (johnweldon1993):

Hmm okay...well start by making a u-sub..... u = x² + 2x + 7 du = ...?

OpenStudy (anonymous):

okay, so du= 2(x+1)dx or..?

OpenStudy (johnweldon1993):

Right! Now look how close that looks to your numerator... divide out the 2 from that and you'll have \[\frac{1}{2}\int\limits_{}^{}\frac{du}{\sqrt{u}}\] or \[\frac{1}{2}\int\limits_{}^{}\frac{1}{\sqrt{u}}du\] What is the integral of 1/√u ?

OpenStudy (isaiah.feynman):

ACTUALLY GUYS,